OpenStudy (anonymous):

Can someone please help me find the lengths of the major and minor axes for: (x-5)^2/4 + (y+3)^2/16 =1

OpenStudy (ranga):

One form of an equation for ellipse is: (x-h)^2 / a^2 + (y-k)^2 / b^2 = 1 where a and b are half of the major/minor axes. Take the bigger one to be the major axis and the smaller one to be the minor axis. compare (x-5)^2/4 + (y+3)^2/16 =1 to (x-h)^2 / a^2 + (y-k)^2 / b^2 = 1 a^2 = 4; so a = 2 b^2 = 16; b = 4 a and b are half the length of the axis. so length of major axis = 2 * b = 2 * 4 = 8 length of minor axis = 2 * a = 2 * 2 = 4